Operations Report – November 22th

Crew 215 Operations Report 22-Nov-2019
SOL: 12
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 12 h 20 min
From what time last night: 1905
To what time this morning: 0725
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 70%
Diesel Reading – 45%
Station Propane Reading – 64%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1460265
Water (static tank) (gal): 258
Static to Loft Pump used – yes
Water in Green Hab (gal): 229.9
Water in Science Dome (gal): 0
Toilet tank emptied: no
Deimos rover used: no, to be sent away for repairs
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: Assigned to Director
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 111.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes
Opportunity rover used: no
Hours: 68
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: no
Hours: 118.1
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: N/A
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used (gal): 0
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With Director
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: N/A
Summary of Hab operations: N/A
Summary of GreenHab operations: Atmospheric equalization to the green hab carried out
during the afternoon, plants were watered (16L,4.6 gal). Door closed at 5pm, heater on
overnight.
Summary of Science Dome operations: Cleaned up lab and mopped floor.
Summary of RAM operations: Wiped benches and vacuumed floor.
Summary of any observatory issues: Nominal
Summary of health and safety issues: New smoke detector in RAM appears to be defective.
Replaced battery twice but continued to sound alarm at odd times.
Questions, concerns and requests to Mission Support: N/A